# Fine-grained Image Retrieval via Dual-Vision Adaptation

### v1

@@ 3.3. Discriminative Perceptual Transfer
(7) $\mathcal{L}_{proxy}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}\right),$
@@
Discriminative Perceptual Transfer. With the discriminative category proxy set C, and the retrieval embeddings $\hat{\textbf{E}}^{R}$ of the origin training images are used for distillation:
(8) $\mathcal{L}_{dpt}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\hat{\textbf{E}}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\hat{\textbf{E}}^{R}\|,\|\textbf{c}\|))}\right).$
3.4. Overall Function
As illustrated in Fig. 2, our objective function comprises two components: the proxy learning loss and the distillation loss $\mathcal{L}_{dpt}$. Finally, the total loss for our framework can be defined as:
(9) $\mathcal{L}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}\right)+\beta\mathcal{L}_{dpt},$
where $\textbf{C}\in\{\textbf{C},\textbf{C}^{\mathrm{b}}\}$ contains original training category and background category, and $\beta$ is the hyper-parameter to weight the $\mathcal{L}_{dpt}$.
4. Experiments
Table 2. Comparison with state-of-the-art methods in the closed-set setting on CUB-200-2011 and Stanford Cars. The best result is shown in bold, and the second-best result is underlined.
@@ 4. Experiments
Method Backbone Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) CNN 80.4 85.7 89.3 92.3 77.2 83.0 87.2 90.2
HIST (Lim et al., 2022) CNN 75.6 83.0 88.3 91.9 89.2 93.4 95.9 97.6
PNCA++  (Teh et al., 2020) CNN 80.4 85.7 89.3 92.3 86.4 92.3 96.0 97.8
Hyp-ViT (Ermolov et al., 2022) ViT 84.2 91.0 94.3 96.0 76.7 85.2 90.8 94.7
DVA (Ours) ViT 85.2 91.4 94.6 96.2 88.0 93.2 97.0 98.5
Table 3. Comparison with state-of-the-art methods in the open-set setting on CUB-200-2011 and Stanford Cars.
@@
Method Backbone Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
PNCA  (Movshovitz-Attias et al., 2017) CNN 49.2 61.9 67.9 72.4 73.2 82.4 86.4 88.7
DGCRL (Zheng et al., 2019) CNN 67.9 79.1 86.2 91.8 75.9 83.9 89.7 94.0
DAS (Liu et al., 2022) CNN 69.2 79.3 87.1 92.6 87.8 93.2 96.0 97.9
IBC (Seidenschwarz et al., 2021) CNN 70.3 80.3 87.6 92.7 88.1 93.3 96.2 98.2
Proxy-Anchor (Kim et al., 2020) CNN 71.1 80.4 87.4 92.5 88.3 93.1 95.7 97.0
HIST (Lim et al., 2022) CNN 71.4 81.1 88.1 - 89.6 93.9 96.4 -
PNCA++  (Teh et al., 2020) CNN 70.1 80.8 88.7 93.6 90.1 94.5 97.0 98.4
FRPT (Wang et al., 2023b) CNN 74.3 83.7 89.8 94.3 91.1 95.1 97.3 98.6
DFML (Wang et al., 2023c) ViT 79.1 86.8 - - 89.5 93.9 - -
Hyp-ViT (Ermolov et al., 2022) ViT 84.0 90.2 94.2 96.4 86.0 91.9 95.2 97.2
DVA (Ours) ViT 84.9 90.6 94.5 96.7 90.7 94.8 97.1 98.4
Table 4. Comparison with state-of-the-art methods on Stanford Dogs.
Open-set Setting Closed-set Setting
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 86.1 91.9 95.7 97.6 83.7 89.9 93.9 96.5
HIST (Lim et al., 2022) 86.2 92.3 95.6 97.5 84.7 90.1 93.4 95.6
PNCA++  (Teh et al., 2020) 86.0 92.3 95.6 97.7 83.8 90.1 94.7 96.8
Hyp-ViT (Ermolov et al., 2022) 87.8 92.8 95.9 97.6 79.2 86.9 91.8 95.2
DVA (Ours) 90.4 94.6 97.0 98.3 87.7 93.0 95.8 97.3
4.1. Experiment Setup
4.1.1. Datasets
CUB-200-2011 (Branson et al., 2014) comprises 11, 788 bird images from 200 bird species. In the closed-set setting, the dataset is divided into training and testing subsets comprising 5,994 and 5,794 images, respectively, out of a total of 11,788 images. For the open-set setting, we employ the first 100 subcategories (comprising 5,864 images) for training, and the remaining subcategories (comprising 5,924 images) are used for testing.
Stanford Cars (Krause et al., 2013) consists of 16,185 images depicting 196 car variants. Similarly, these images were split into 8,144 training images and 8,041 test images in the closed-set setting. For the open-set setting, we utilize the first 98 subcategories (comprising 8,054 images) for training and the remaining 98 subcategories (comprising 8,131 images) for testing.
Stanford Dogs (Dataset, 2011) contains 20,580 images showcasing dogs across 120 subcategories. We use the first 60 subcategories with 12,000 images for training and the remaining 60 subcategories with 8580 images for testing for the closed-set setting. For the open-set setting, we utilize the first 60 subcategories (comprising 10,651 images) for training and the remaining 60 subcategories (comprising 9,929 images) for testing.
NABirds (Horn et al., 2015) contains 48,562 images showcasing North American birds across 555 subcategories. We use the entire dataset as the generalization test dataset.
FGVC Aircraft (Maji et al., 2013) contains 10,000 images showcasing 100 subcategories. We use the entire dataset as the generalization test.
INat2021 (Van Horn et al., 2021) contains more than 2.7 million images of 10,000 categories. For the convenience of evaluation, we use the 27,105 images from the first 100 subcategories in its training set as the generalization test dataset.
4.1.2. Evaluation protocols
To evaluate retrieval performance, we adopt Recall@K with cosine distance in previous work (Song et al., 2016), which calculates the recall scores of all query images in the test set. For each query image, the top K similar images are returned. A recall score of 1 is assigned if at least one positive image among the top K images; otherwise, it is 0.
4.1.3. Implementation details
In experiments, we employ the ViT-B-16 (Dosovitskiy et al., 2021) pre-trained on ImageNet21K (Russakovsky et al., 2015) as our image encoder. All input images are resized to 256 $\times$ 256, and crop them into 224 $\times$ 224. In the training stage, we utilize the Adam optimizer with weight decay of 1e-4, and employ cosine annealing as the optimization scheduler. The learning rate for all datasets is initialized to 1e-1 except for Stanford Dogs, which has a learning rate of 1e-2. The number of training epochs for all datasets is set to 10 and the batch size is set to 32. We train our model on a single NVIDIA 3090 GPU to accelerate the training process.
4.2. Comparison with State-of-the-Art Methods
4.2.1. Closed-set Setting
We first compare our proposed DVA with previous competitive methods under closed-set setting. Quantitative comparison results for the CUB-200-2011 and Stanford Cars datasets are presented in Table 2, and the results for the Stanford Dogs dataset can be found in Table 4. From these tables, it can be observed that our proposed DVA outperforms other state-of-the-art methods on CUB-200-2011 and Stanford Dogs, and achieves competitive performance on Stanford Cars, which demonstrates the enhanced discriminative capability of our DVA for fine-grained visual retrieval. Specifically, in comparison with Hyp-ViT (Ermolov et al., 2022), the current state-of-the-art on CUB-200-2011, our DVA demonstrates a 1.0% improvement in Recall@1. The experimental results on the Stanford Cars indicate that our method outperforms the most of existing methods but falls slightly behind HIST (Lim et al., 2022). We argue the possible reason may be attributed to the relatively regular and simpler shape of the cars.
4.2.2. Open-set Setting
The open-set setting poses greater challenges compared to the closed-set setting due to the unknown subcategories in the test data. The experimental results for the CUB-200-2011 and Stanford Cars datasets are shown in Table 3, while the results for the Stanford Dogs dataset are provided in Table 4. The results reveal a consistent trend in both open and closed-set settings: our method outperforms other state-of-the-art approaches on CUB-200-2011 and Stanford Dogs, while delivering competitive results on the Stanford Cars dataset. To be specific, in comparison with Hyp-ViT (Ermolov et al., 2022) on CUB-200-2011, our DVA exhibits a 0.9% improvement in Recall@1. Experimental results on the Stanford Cars dataset show that our method outperforms most existing methods but lags behind FRPT (Wang et al., 2023b) by a slight margin. The results of experiments on Stanford Dogs are shown in Table 4. Our DVA demonstrates a clear performance advantage over Hyp-ViT, further indicating its superior generalizability.
4.3. Generalizability Analysis
To further demonstrate the generalization ability of our method, we conduct comparative experiments on the out-of-distribution datasets NABirds (Horn et al., 2015), FGVC Aircraft (Maji et al., 2013), and INat2021 (Van Horn et al., 2021). The results are presented in Tables 5, 6, and 7, respectively. Our method consistently outperforms all other approaches across these datasets. Notably, it demonstrates a clearer advantage compared to the results discussed in Section 4.2, highlighting its strong generalization capability. These findings confirm that DVA not only distinguishes similar samples within the distribution effectively, but also performs well on similar samples outside the distribution. This robust generalization ability is crucial for real-world applications where data distributions can vary significantly.
Table 5. Comparison on the NABirds dataset with state-of-the-art methods trained on the CUB-200-2011 and Stanford Cars datasets.
CUB-200-2011 Standford Cars
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 51.0 61.3 70.3 77.9 19.6 26.5 34.6 44.0
HIST (Lim et al., 2022) 50.4 60.3 69.3 76.8 14.4 20.1 26.4 34.4
PNCA++  (Teh et al., 2020) 57.9 68.7 77.5 84.7 36.5 46.5 56.9 66.9
Hyp-ViT (Ermolov et al., 2022) 64.7 74.6 82.7 88.7 44.6 55.3 65.3 74.3
DVA (Ours) 79.2 86.4 91.5 94.8 70.5 80.1 87.2 92.1
Table 6. Comparison on the FGVC Aircraft dataset with state-of-the-art methods trained on the CUB-200-2011 and Stanford Cars datasets.
CUB-200-2011 Standford Cars
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 28.7 37.6 46.9 57.0 32.3 41.4 51.3 62.0
HIST (Lim et al., 2022) 31.4 39.9 49.3 59.3 32.4 41.3 50.6 60.8
PNCA++  (Teh et al., 2020) 38.9 48.8 58.8 69.2 36.9 47.1 57.3 67.8
Hyp-ViT (Ermolov et al., 2022) 39.3 49.3 59.3 69.4 42.6 52.2 62.5 72.0
DVA (Ours) 42.2 52.1 62.2 72.3 48.3 56.2 68.1 77.2
Table 7. Comparison on the INat2021 dataset with state-of-the-art methods trained on the CUB-200-2011 and Stanford Cars datasets.
CUB-200-2011 Standford Cars
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 24.0 33.4 44.0 55.4 13.7 20.9 30.1 41.5
HIST (Lim et al., 2022) 21.1 29.9 40.5 52.0 13.4 20.1 29.2 40.5
PNCA++  (Teh et al., 2020) 30.1 40.7 52.0 64.0 25.6 35.5 47.3 59.4
Hyp-ViT (Ermolov et al., 2022) 38.4 50.2 62.1 72.6 31.7 43.0 55.1 66.9
DVA (Ours) 51.8 63.7 73.9 82.3 45.0 57.2 68.3 78.2
4.4. Ablation Studies and Analysis
Table 8. The Recall@K results (%) of component ablation study on CUB-200-2011.
Setting R@1 R@2 R@4 Latency (ms)
Base 72.5 82.6 89.1 1.51
Base + ICA 83.6 90.3 93.9 1.51
Base + ICA + OPA 86.8 92.0 94.7 3.65
Base + ICA + OPA + DPT 84.9 90.6 94.5 1.51
4.4.1. Efficacy of various components
The proposed DVA comprises three essential components: Object-Perceptual Adaption (OPA), In-Context Adaptation (ICA), and Discriminative Perceptual Transfer (DPT). We conducted ablation experiments on these components, and the results are reported in Table 8, with the Base representing the pure ViT. The introduction of ICA improves the Recall@1 accuracy by 11.1% on CUB-200-2011, which shows that ICA effectively guides the FGIR task to the task solved in the original process through feature adjustment. OPA further improves the ability of the pre-trained model to distinguish similar subcategories on FGIR, bringing a gain of 3.2% Recall@1 on CUB-200-2011, but it brings huge computational overhead. DPT effectively removes the huge computational overhead brought by OPA and achieves a balance between performance and retrieval efficiency.
4.4.2. Sensitivity analysis of parameter $\beta$
Figure 3. Analyses of hyper-parameter $\beta$ on CUB-200-2011.
As illustrated in Fig. 3, we investigate the impact of varying values of $\beta$, representing the weight of loss in Eq. (9). The performance gradually increases with the value but decreases when the value exceeds 3. The drop in performance can be attributed to the fact that as the value of $\beta$ increases, the learned feature space tends to receive samples adjusted by the OPA module, while the input at inference is not processed by the OPA, resulting in a difference in the input space that causes performance degradation. Consequently, we set $\beta$ to 3 for all datasets.
Figure 4. Examples of top-10 retrieved images on CUB-200-2011 by DVA. The images with green boxes are the correct ones, and those with red boxes are the wrong ones.
Table 9. The Recall@K results (%) of component ablation study on CUB-200-2011.
Setting R@1 R@2 R@4 R@8 Prams (M)
$\textbf{IC}_{q}$ 82.7 89.1 92.9 95.6 0.295
$\textbf{IC}_{k}$ 81.8 88.8 93.1 95.5 0.295
$\textbf{IC}_{v}$ 84.0 90.0 93.6 95.6 0.295
$\textbf{IC}_{q}$, $\textbf{IC}_{k}$ 84.9 90.6 94.5 96.7 0.590
$\textbf{IC}_{q}$, $\textbf{IC}_{v}$ 84.5 90.5 93.9 95.7 0.590
$\textbf{IC}_{k}$, $\textbf{IC}_{v}$ 82.6 89.0 93.4 95.4 0.590
$\textbf{IC}_{q}$, $\textbf{IC}_{k}$, $\textbf{IC}_{v}$ 84.4 90.2 93.8 95.9 0.885
4.4.3. Analysis of In-Context Adaptation
To validate the effectiveness of the proposed ICA components, we further divide the ICA module into three parts: $\textbf{IC}_{q}$, $\textbf{IC}_{k}$, and $\textbf{IC}_{v}$, corresponding to the Q, K, and V projectors in the Attention Projector, respectively. We perform ablation studies on these components, and the results are presented in Table 9. The results show that using $\textbf{IC}_{q}$ alone or in combination with $\textbf{IC}_{k}$ and $\textbf{IC}_{v}$ leads to better performance. Notably, the best results are achieved when $\textbf{IC}_{q}$ and $\textbf{IC}_{k}$ are used together. Therefore, we adopt this combination to construct the final ICA module in our DVA.
4.4.4. Visualization
After visualizing the top 10 retrieval results on CUB-200-2011 in Fig. 4, we observe that DVA effectively retrieves images from the same subcategory, even across different variations and backgrounds. This demonstrates that our dual-vision adaptation mechanism enhances the pre-trained model’s ability to focus on fine-grained characteristics. For the failure case in the second row, even humans can easily overlook the subtle differences between the query image and the returned image and make mistakes. To explore the superiority of the designed DVA, We also conduct visualization as shown in Fig. 5. The first column displays the input images. The second column shows the class activation maps generated by GradCAM for the baseline model, which processes the input images. The third column presents the class activation maps for DVA using the same input. The results indicate that DVA effectively reduces attention to background regions and improves the baseline model’s focus on critical areas, capturing finer details in subcategories such as the head, wings, and tail.
Figure 5. Class activation visualizations on CUB-200-2011. For each sample, we show the input image, the class activation map of the baseline model, and the proposed DVA.
4.4.5. Comparison with other efficient parameter fine-tuning methods
We compare our method with two efficient parameter fine-tuning approaches: CLIP-Adapter (Gao et al., 2024) and AdaptFormer (Chen et al., 2022a). As shown in Table 10, both methods perform well on the CUB-200-2011 dataset. However, our DVA still achieves superior results, demonstrating its ability to capture subtle differences and accurately distinguish similar subcategories through dual-vision adaptation. Furthermore, DVA outperforms the fully fine-tuned ViT. This is expected, as fine-tuning a pre-trained model on a small fine-grained dataset often causes overfitting, reducing its generalization capability.
Table 10. The Recall@K results (%) of different parameter efficient fine-tuning methods on CUB-200-2011.
Methods R@1 R@2 R@4 R@8 Params (M)
Fine-tuning ViT 82.7 88.7 92.5 95.2 85.6
CLIP-Adapter (Gao et al., 2024) 69.8 80.0 87.5 92.1 0.295
AdaptFormer (Chen et al., 2022a) 73.3 83.4 89.3 93.2 0.590
DVA (Ours) 84.9 90.6 94.5 96.7 0.590
5. Conclusion
In this paper, we propose Dual-Vision Adaptation (DVA), a method aimed at improving generalization in fine-grained image retrieval. While conventional methods risk overfitting the training set due to full-model fine-tuning, DVA mitigates this issue by preserving pre-trained knowledge. DVA introduces Object-Perceptual Adaption (OPA) and In-Context Adaptation (ICA) to guide a frozen pre-trained vision model in performing fine-grained retrieval tasks. Specifically, OPA leverages a visual foundation model to create a new background category and enhance pixels contributing to category prediction, and helping the model better perceive critical objects and their distinguishing features. ICA introduce a small set of in-context adaptation parameters that enables the frozen pre-trained model to capture discriminative details through feature adaptation. Additionally, DVA achieves a balance between discriminative perception and retrieval efficiency with the design of Discriminative Perceptual Transfer (DPT). Extensive experiments on three in-distribution and three out-of-distribution fine-grained benchmarks show that our DVA achieves excellent performance with very few learnable parameters.

### v4

@@ Introduction
$\begin{split}\hat{\mathcal{L}}_{proxy}&=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\hat{\textbf{C}}}\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}\right),\\ \mathcal{L}&=\hat{\mathcal{L}}_{proxy}+\beta\mathcal{L}_{dpt},\end{split}$ (5)
@@
Experiments
Experiment Setup
Datasets.
CUB-200-2011 (cub) comprises 11, 788 bird images from 200 bird species. In the closed-set setting, the dataset is divided into training and testing subsets comprising 5,994 and 5,794 images, respectively, out of a total of 11,788 images. For the open-set setting, we employ the first 100 subcategories (comprising 5,864 images) for training, and the remaining subcategories (comprising 5,924 images) are used for testing.
Stanford Cars (stanfordcar) consists of 16,185 images depicting 196 car variants. Similarly, these images were split into 8,144 training images and 8,041 test images in the closed-set setting. For the open-set setting, we utilize the first 98 subcategories (comprising 8,054 images) for training and the remaining 98 subcategories (comprising 8,131 images) for testing.
Stanford Dogs (dataset2011novel) contains 20,580 images showcasing dogs across 120 subcategories. We use the 120 subcategories with 12,000 images for training and the remaining 120 subcategories with 8580 images for testing for the closed-set setting. For the open-set setting, we utilize the first 60 subcategories (comprising 10,651 images) for training and the remaining 60 subcategories (comprising 9,929 images) for testing.
Evaluation protocols.
To evaluate retrieval performance, we adopt Recall@K with cosine distance in previous work (recallk), which calculates the recall scores of all query images in the test set. For each query image, the top K similar images are returned. A recall score of 1 is assigned if at least one positive image among the top K images; otherwise, it is 0.
Implementation details.
In experiments, we employ the ViT-B-16 (vit) pre-trained on ImageNet21K (imagenet) as our image encoder. All input images are resized to 256 $\times$ 256, and crop them into 224 $\times$ 224. In the training stage, we utilize the Adam optimizer with weight decay of 1e-4, and employ cosine annealing as the optimization scheduler. The learning rate for all datasets is initialized to 1e-1 except for Stanford Dogs, which has a learning rate of 1e-2. The number of training epochs for all datasets is set to 10 and the batch size is set to 32. We train our model on a single NVIDIA 3090 GPU to accelerate the training process.
Comparison with State-of-the-Art Methods
@@
Method R@1 R@2 R@4 R@8 R@1 R@2 R@4 R@8
DAS (das) 69.2 79.3 87.1 92.6 87.8 93.2 96.0 97.9
IBC (ibc) 70.3 80.3 87.6 92.7 88.1 93.3 96.2 98.2
Proxy-Anchor (proxy) 71.1 80.4 87.4 92.5 88.3 93.1 95.7 97.0
HIST (hist) 71.4 81.1 88.1 - 89.6 93.9 96.4 -
PNCA++ (pnca++) 70.1 80.8 88.7 93.6 90.1 94.5 97.0 98.4
FRPT (frpt) 74.3 83.7 89.8 94.3 91.1 95.1 97.3 98.6
DFML (dfml) 79.1 86.8 - - 89.5 93.9 - -
Hyp-ViT (Hyp-ViT) 84.0 90.2 94.2 96.4 86.0 91.9 95.2 97.2
DVA (Ours) 84.9 90.6 94.5 96.7 90.7 94.8 97.1 98.4
Table 3: Comparison with state-of-the-art methods in the open-set setting on CUB-200-2011 and Stanford Cars.
Open-set Setting Closed-set Setting
Method R@1 R@2 R@4 R@8 R@1 R@2 R@4 R@8
Proxy-Anchor (proxy) 86.1 91.9 95.7 97.6 83.7 89.9 93.9 96.5
HIST (hist) 86.2 92.3 95.6 97.5 84.7 90.1 93.4 95.6
PNCA++  (pnca++) 86.0 92.3 95.6 97.7 83.8 90.1 94.7 96.8
Hyp-ViT (Hyp-ViT) 87.8 92.8 95.9 97.6 79.2 86.9 91.8 95.2
DVA (Ours) 90.4 94.6 97.0 98.3 87.7 93.0 95.8 97.3
Table 4: Comparison with state-of-the-art methods on Stanford Dogs.
Closed-set Setting.
We first compare our proposed DVA with previous competitive methods under closed-set setting. Quantitative comparison results for the CUB-200-2011 and Stanford Cars datasets are presented in Table 2, and the results for the Stanford Dogs dataset can be found in Table 4. From these tables, it can be observed that our proposed DVA outperforms other state-of-the-art methods on CUB-200-2011 and Stanford Dogs, and achieves competitive performance on Stanford Cars, which demonstrates the enhanced discriminative capability of our DVA for fine-grained visual retrieval. Specifically, in comparison with Hyp-ViT (Hyp-ViT), the current state-of-the-art on CUB-200-2011, our DVA demonstrates a 1.0% improvement in Recall@1. The experimental results on the Stanford Cars indicate that our method outperforms the most of existing methods but falls slightly behind HIST (hist). We argue the possible reason may be attributed to the relatively regular and simpler shape of the cars.
Open-set Setting.
The open-set setting poses greater challenges compared to the closed-set setting due to the unknown subcategories in the test data. The experimental results for the CUB-200-2011 and Stanford Cars datasets are shown in Table 3, while the results for the Stanford Dogs dataset are provided in Table 4. The results reveal a consistent trend in both open and closed-set settings: our method outperforms other state-of-the-art approaches on CUB-200-2011 and Stanford Dogs, while delivering competitive results on the Stanford Cars dataset. To be specific, in comparison with Hyp-ViT (Hyp-ViT) on CUB-200-2011, our DVA exhibits a 0.9% improvement in Recall@1. Experimental results on the Stanford Cars dataset show that our method outperforms most existing methods but lags behind FRPT (frpt) by a slight margin. We attribute FRPT’s performance gains to its carefully designed but computationally expensive modules. In contrast, our method captures subtle differences through dual-vision adaptation while maintaining computational efficiency. The results of experiments on Stanford Dogs are shown in Table 4. Our DVA demonstrates a clear performance advantage over Hyp-ViT, further indicating its superior generalizability.
Ablation Studies and Analysis
Setting R@1 R@2 Latency (ms)
Base 72.5 82.6 1.51
Base + ICA 83.6 90.3 1.51
Base + ICA + OPA 86.8 92.0 3.65
Base + ICA + DP + DPT 84.1 90.2 1.51
Base + ICA + OPA + DPT 84.9 90.6 1.51
Table 5: The Recall@K results (%) of component ablation study on CUB-200-2011.
Efficacy of various components. The proposed DVA comprises three essential components: Object-Perceptual Adaption (OPA), In-Context Adaptation (ICA), Discriminative Perception (DP) of the OPA, and Discriminative Perceptual Transfer (DPT). We conducted ablation experiments on these components, and the results are reported in Table 5, with the Base representing the pure ViT. The introduction of ICA improves the Recall@1 accuracy by 11.1% on CUB-200-2011, which shows that ICA effectively guides the FGIR task to the task solved in the original process through feature adaptation. OPA further improves the ability of the pre-trained model to distinguish similar subcategories on FGIR, bringing a gain of 3.2% Recall@1 on CUB-200-2011, but it brings huge computational overhead. DPT effectively removes the huge computational overhead brought by OPA and achieves a balance between performance and retrieval efficiency. In particular, when only DP is used but no Object Perception, the performance decreases. This shows that Object Perception effectively helps DVA perceive objects and thus better distinguish similar objects.
@@
Setting Recall@1 Recall@2 Recall@1 Recall@2 Prams (M)
$\textbf{IC}_{q}$ 82.7 89.1 87.2 92.9 0.295
$\textbf{IC}_{k}$ 81.8 88.8 85.6 91.7 0.295
$\textbf{IC}_{v}$ 84.0 90.0 86.4 92.3 0.295
$\textbf{IC}_{q}$, $\textbf{IC}_{k}$ 84.9 90.6 90.7 94.8 0.590
$\textbf{IC}_{q}$, $\textbf{IC}_{v}$ 84.5 90.5 90.3 94.5 0.590
$\textbf{IC}_{k}$, $\textbf{IC}_{v}$ 82.6 89.0 89.5 94.1 0.590
$\textbf{IC}_{q}$, $\textbf{IC}_{k}$, $\textbf{IC}_{v}$ 84.4 90.2 90.4 94.6 0.885
Table 6: The Recall@K results (%) of component ablation study on CUB-200-2011 and Stanford Cars.
Analysis of In-Context Adaptation. To validate the effectiveness of the proposed ICA components, we further divide the ICA module into three parts: $\textbf{IC}_{q}$, $\textbf{IC}_{k}$, and $\textbf{IC}_{v}$, corresponding to the Q, K, and V projectors in the Attention Projector, respectively. We perform ablation studies on these components, and the results are presented in Table 6. The results show that using $\textbf{IC}_{q}$ alone or in combination with $\textbf{IC}_{k}$ and $\textbf{IC}_{v}$ leads to better performance. Notably, the best results are achieved when $\textbf{IC}_{q}$ and $\textbf{IC}_{k}$ are used together. Therefore, we adopt this combination to construct the final ICA module in our DVA.
Sensitivity analysis of parameter $\beta$. As illustrated in Fig. 3, we investigate the impact of varying values of $\beta$ in Eq. (5). The performance gradually increases with the value but decreases when the value exceeds 3. The drop in performance can be attributed to the fact that as the value of $\beta$ increases, the learned feature space tends to receive samples adapted by the OPA module, while the input at inference is not processed by the OPA, resulting in a difference in the input space that causes performance degradation. Consequently, we set $\beta$ to 3 for all datasets.
Figure 3: Analyses of hyper-parameter $\beta$ on CUB-200-2011.
Visualization.
To explore the superiority of the designed DVA, We also conduct visualization as shown in Fig. 4. The first column displays the input images. The second column shows the class activation maps generated by GradCAM for the baseline model, which processes the input images. The third column presents the class activation maps for DVA using the same input. The results indicate that DVA effectively reduces attention to background regions and improves the baseline model’s focus on critical areas, capturing finer details in subcategories such as the head, wings, and tail.
Figure 4: Class activation visualizations on CUB-200-2011. For each sample, we show the input image, the class activation map of the baseline model, and the proposed DVA.
Conclusion
In this paper, we present Dual Visual Adaptation (DVA), a novel approach designed to alleviate overfitting in fine-grained image retrieval by retaining knowledge from pre-training. DVA integrates three key components: Object-Perceptual Adaptation, which introduces a new background category and enhances discriminative object regions using a visual foundation model; In-Context Adaptation, which employs lightweight parameters to adapt fine-grained features within a frozen backbone; and Discriminative Perceptual Transfer, which distills discriminative knowledge into the encoder for efficient inference. Extensive experiments on three fine-grained benchmarks demonstrate that DVA achieves strong performance with simple yet effective dual-vision adaptation.